OpenStudy (anonymous):

on which of the following lines does the point (-2 , -3) lie? A.y = 4x + 4 b.y = 3x - 8 c. y = -2x + 1 d. y = -4x - 4

OpenStudy (anonymous):

help plz so i can pass!!! ill give you a metal metal and ill become a fan!

OpenStudy (jack1):

try each y = 4x + 4, if x = -2 and y = -3 -3 = 4(-2) + 4 -3 = -8 + 4 -3 = -4 LHS doesn't = RHS, so not A rinse and repeat bro

OpenStudy (midhun.madhu1987):

The point (-2,-3) indicates that x=-2 and y=-3. Just substitute the values of x and y in all the equations and check which of them satisfies the equality...

OpenStudy (anonymous):

so its c? help plz anyone

OpenStudy (mathmath333):

is the question correct ?

OpenStudy (anonymous):

yes is it? anyone help plz :(

OpenStudy (midhun.madhu1987):

i think the options are wrong...

OpenStudy (mathmath333):

none of the 4 options match

OpenStudy (midhun.madhu1987):

if option c is y=2x+1, then it will be correct

OpenStudy (mathmath333):

check it for sure

OpenStudy (anonymous):

a. y = 4x + 4 b. y = 3x - 8 c. y = -2x + 1 d. y = -4x - 4

OpenStudy (anonymous):

guys give me an answer please! i only have 2 mins left to answer!!!! ill give you a metal and ill become a fan!